3.) For this problem you need to generate data so open the probability simulator on the TI calculator and set your seed (Use3.) For this problem you need to generate data so open the probability simulator on the TI calculator and set your seed (UseSEED NUMBER: 3932
Answer #1

obsa ho X KN (x1-7)- 7 2 25 2 10 2 25 3 7 3 16 3 ज 5 4 S ।। G 6 ।। 7 6 7 8 7 88 7 92 7 107 B ।) 8 B 12-2 8 13 7 10 ما । 4 10an The sample mean V. Eri 112 16 IT b) The sample standard deviation = Ecxi-x)? h- ** x 124 20.6667 4-5460 c) The median 194 th. 4 thobsn +0.25 (5th 0.75( 13th d) To find Igr. IOR: 02-01 9 1 = (n+13th obsh (16+ )*hob seruation = (4.25) thobisin += 8 + 0.75(10.5) st 0.75 (2) 9.5 I GR 93 91 9.55.25 4.25 CSS Scanned with CamScanner